OpenStudy (anonymous):

what is the radius of a circle is the circumference is 56 pi cm squared

OpenStudy (campbell_st):

well the formula for circumference in terms of radius is \[C = 2 \times \pi \times r\] using your information you know \[56\pi = 2 \times \pi \times r\] now solve for r